Two different rectangles each have an area of 360 sq. centimeters. The length of the second rectangle is 12 centimeters greater than the first, whereas its width is 5 centimeters less than the first rectangle. Find the difference of the perimeters of the two rectangles.
Answer by uma(370)

Let the length of the first rectangle be = x cm
So length of the second rectangle = (x+12) cm
Let the width of the first rectangle = y cm
So width of the second rectangle = (y-5) cm
perimeter of the first rectangle = 2(x + y)
= 2x + 2y cm
Perimeter of the second rectangle = 2[(x+12) + (y-5)]
= 2[x+12+y-5]
= 2(x + y + 7)
= 2x + 2y + 14 cm
Difference in their perimeters = 2x + 2y + 14 - (2x + 2y)
= 2x + 2y + 14 - 2x - 2y
= 14 cm

Good Luck!!!